OpenStudy (goose.):

Find the solution of 4 times the square root of the quantity of x plus 2 equals negative 16, and determine if it is an extraneous solution. x = 14; not extraneous x = 2; extraneous x = 2; not extraneous x = 14; extraneous

OpenStudy (goose.):

\[4\sqrt{x+2}=-16\]

OpenStudy (anonymous):

any ideas on how to start?

OpenStudy (goose.):

You would move -16 to the left side, and set it equal to 0?

OpenStudy (anonymous):

I'd divide by the 4.

OpenStudy (anonymous):

It's not a polynomial, so you don't have to set it equal to 0

OpenStudy (goose.):

So I'd have? \[\frac{ \sqrt{x+2} }{ 4 }=-4\]

OpenStudy (anonymous):